OpenStudy (anonymous):

How would the expression (x^2+1)(y^2+4) be rewritten using Two Squares? A. (xy-2)^2+(x+2y)^2 B. (xy-2)^2+(2x+y)^2 C. (xy-2)^2-(4x+y)^2 D.(xy+2)^2+(x+2y)^2

OpenStudy (trojanpoem):

y^2 x^2 + 4x^2 + y^2 + 4 (yx +2)^2 + (2x+y)^2

OpenStudy (anonymous):

what

OpenStudy (loser66):

It doesn't make sense to you, right? and "How can I know that way", right? Suggestion: just open parentheses the options (ONE BY ONE), and simplify them to see which one gives you the given expression. That way doesn't ask you the experience or guessing.

OpenStudy (loser66):

I do A for you as example A) \((xy-2)^2 +(x+2y)^2\\=(x^2y^2-4xy+4)+(x^2+4xy+4y^2)\\=x^2+4y^2+x^2y^2+4\) while your give expression is \((x^2+1)(y^2+4)=x^2y^2+4x^2+y^2+4\) They are different, right? hence A is NOT the right answer.

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

its B

OpenStudy (loser66):

I gave you the way to do. You do it carefully and take responsibility for your work. I don't check.

OpenStudy (anonymous):

I didn't say you did?

OpenStudy (loser66):

No, you didn't. Just gave the reason why I didn't say Yes or No for your answer.

OpenStudy (anonymous):

ok